OpenStudy (anonymous):

How to determine if a function is even, odd, or neither? y = -6x^3 - 4x I'd say that the function is odd because of the odd exponent but I'm not sure because x is being multiplied and not just cubed.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a polynomial with all odd exponents is odd hence the name (although that is not the definition)

OpenStudy (yamyam70):

You may be asked to "determine algebraically" whether a function is even or odd. To do this, you take the function and plug –x in for x, and then simplify. If you end up with the exact same function that you started with (that is, if f(–x) = f(x), so all of the signs are the same), then the function is even. If you end up with the exact opposite of what you started with (that is, if f(–x) = –f(x), so all of the "plus" signs become "minus" signs, and vice versa), then the function is odd. In all other cases, the function is "neither even nor odd".

OpenStudy (yamyam70):

source : http://www.purplemath.com/modules/fcnnot3.htm

OpenStudy (anonymous):

your concern about the coefficients is a good one, but they make no difference in this case

OpenStudy (anonymous):

oh i totally disagree

OpenStudy (agent0smith):

^ for polynomials.

OpenStudy (agent0smith):

Disagree with...?

OpenStudy (anonymous):

Last question is this: I'm supposed to write the equation that would make the parent function (solid line) transform to the related function (dashed line), here's a picture of the graph: http://puu.sh/4Kx7R.png I wrote f(x) = |x^3 + 3| + 4 I'm not sure if it should be considered an absolute value or if it should, but I know that +3 is right because it moves to the left. But, since +4 is outside the absolute value, I'm not sure if it should be -4 even though it moves up.

OpenStudy (anonymous):

question is "is the function odd" \[f(x)=x^3+6x\] is odd, \[f(x)=x^3+6x^2\] is not

OpenStudy (anonymous):

for your second question, there is no absolute value at all, get rid of it and replace it by parentheses

OpenStudy (agent0smith):

Oh i might actually be mixing this up with just the end behaviour of polynomials...

OpenStudy (anonymous):

okay, \[f(x) = (x^3 + 3) + 4\] ?

OpenStudy (anonymous):

that looks much better

OpenStudy (anonymous):

oh no hold the phone

OpenStudy (agent0smith):

(x+3)^3 + 4

OpenStudy (anonymous):

\[f(x)=(x+3)^3+4\] moves left 3, up 4

OpenStudy (anonymous):

Gotcha. Thank you!

OpenStudy (agent0smith):

we have the numbers wrong anyway... isn't it only left 2, up 3? hard to tell on that tiny graph.

OpenStudy (anonymous):

yeah it is hard to tell for sure i would go with 3, but i could be wrong

OpenStudy (anonymous):

I got 3 left and 4 up.

OpenStudy (anonymous):

now i go with 2 look at the part that looks almost horizontal that is two units to the left of the origin

OpenStudy (agent0smith):

And you only need to write it in terms of the parent function. Not actually write the function. If f(x) is your parent function, f(x+2) + 3 is your transformed one.

OpenStudy (anonymous):

the part of \(f(x)=x^3\) that is flat looking is at \((0,0)\) and the part of the other one that is flat looking is at \((-2,3)\)

OpenStudy (anonymous):

oh crap i think it is 2

OpenStudy (anonymous):

yeah me too, and up 3 rather than 4

OpenStudy (anonymous):

aw man. :(

OpenStudy (anonymous):

on line class?

OpenStudy (anonymous):

yup, and already sent it haha.. i'm pretty sure it's fine though, i'm confident in the rest of my answers so it's cool.

OpenStudy (anonymous):

whew!

OpenStudy (anonymous):

If anything I can email my teacher tonight and tell her that I think I messed it up, maybe she'll give me credit for it. Thanks guys!

OpenStudy (agent0smith):

Tell her it's her fault for giving such tiny graphs :P

OpenStudy (agent0smith):

And sorry @meegan if i confused you with my post earlier... i deleted it because i was thinking of end behaviour of odd-highest power vs even-highest power polynomials... recently taught that to my precalc class and I saw the odd and even and didn't really think it through :P